OpenStudy (anonymous):

Choose the equivalent system of linear equations that will produce the same solution as the one given below. 4x - y = -11 2x + 3y = 5 -4x - 9y = -19 -10y = -30 4x + 3y = 5 2y = -6 7x - 3y = -11 9x = -6 12x - 3y = -33 14x = -28

OpenStudy (anonymous):

the top equation is the one youre supposed to base the answer off of sorry if that was not clear the rest are multiple choice answers

OpenStudy (texaschic101):

4x - y = -11 ---> y = 4x + 11 2x + 3y = 5 2x + 3(4x + 11) = 5 2x + 12x + 33 = 5 2x + 12x = 5 - 33 14x = - 28 x = -2 4x - y = -11 4(-2) - y = -11 -8 - y = -11 -y = -11 + 8 -y = -3 y = 3 so your solution should be (-2,3) now we have to find the one with the same solution by subbing in our answer into each answer choice to see which one comes out true. Lets try this one.. -4x - 9y = -19 -4(-2) - 9(3) = -19 8 - 27 = -19 -19 = -19 (correct) -10y = -30 -10(3) = -30 -30 = -30 (correct) answer is first answer choice

OpenStudy (anonymous):

thankyou so much!

OpenStudy (anonymous):

would you mind helping me with one more? \